OpenStudy (anonymous):

use l'hospital rule for this limits 1) lim(1+2/x)^3x approch x to infinity please respond to me

OpenStudy (sirm3d):

L'HR cannot be used immediately. first, let \[y=\lim_{x \rightarrow +\infty} \left(1+\frac{2}{x}\right)^{3x}\]

OpenStudy (anonymous):

okay then i will use ln for both side

OpenStudy (sirm3d):

right.

OpenStudy (anonymous):

then i don't know to complete please could you solve it

OpenStudy (sirm3d):

i'll do some parts. \[\ln y=\ln \lim_{x \rightarrow +\infty}\left(1+\frac{2}{x}\right)^{3x}= \lim_{x \rightarrow +\infty}\ln \left(1+\frac{2}{x}\right)^{3x}\]

OpenStudy (anonymous):

2\x

OpenStudy (anonymous):

and the lim xrightarrow infty

OpenStudy (sirm3d):

apply the property of the \(\ln\) function,\[\ln y=\lim_{x \rightarrow +\infty} 3x \ln \left(1+\frac{2}{x}\right)\]and rewrite as \[\ln y=3\lim_{x \rightarrow +\infty} \frac {\ln \left(1+\frac{2}{x}\right)}{\frac{1}{x}}\]

OpenStudy (sirm3d):

now, it's your turn to apply the L'HR since the rational expression is an indeterminate form \[\left[ \frac{0}{0}\right]\]

OpenStudy (anonymous):

okay thank you

OpenStudy (raden):

actually, it will be easier use the define of